VIBRATIONAL SPECTRA OF A HEXA-ATOMIC SILICON-CARBON CLUSTER: LINEAR SiC_{4}Si ^{a}

\underline{J. D. PRESILLA-MÁRQUEZ}, C. M. L. RITTBY, AND W. R. M. GRAHAM, Department of Physics, Texas Christian University, Fort Worth, TX 76129.

Ongoing Fourier transform infrared investigations of the spectra and structures of silicon-carbon clusters trapped in Ar at 10 K, have resulted in the observation of the first six-member cluster, Si_{2}C_{4} cluster. Two vibrational fundamentals of the linear SiC_{4}Si structure have been assigned, the C=C stretching mode ν_{4}(σ_{u}) = 1807.4 cm^{−1}, and the Si-C stretching mode ν_{5}(σ_{u}) = 719.1 cm^{−1}. The measured frequencies, relative intensities, and ^{13}C and ^{29,30}Si isotopic shifts are in very good agreement with the predictions of DFT calculations and confirm the previously predicted, linear symmetric geometry for the ground state of SiC_{4}Si.

aResearch supported by the Welch Foundation